# FINITE SUBGROUPS OF MAXIMAL ORDER OF THE CREMONA GROUP OVER THE RATIONALS

### v1

@@ 4. Del Pezzo surfaces of degree $6$
Over the algebraic closure of $\mathbb{Q}$, we have an isomorphism $S_{\overline{\mathbb{Q}}}\rightarrow X_{\overline{\mathbb{Q}}}$. Let $L/\mathbb{Q}$ be the smallest Galois field extension such that the six $(-1)-$curves are defined over $L$. The surface $S$ has six $(-1)-$curves defined over $L$ but not necessarily over $\mathbb{Q}$. The conjugation of the action gives an action of $\operatorname{Gal}(L/\mathbb{Q})$ on $S_{L}$. Indeed, if we write the isomorphism $\varphi:S_{L}\rightarrow X_{L}$ and let $g\in\operatorname{Gal}(L/\mathbb{Q})$, then the following diagram
$\begin{array}[]{ccc}\hskip 14.22636ptS_{L}&\stackrel{{\scriptstyle\varphi}}{{\rightarrow}}&X_{L}\\ g\downarrow&&\downarrow\tilde{g}\\ \hskip 14.22636ptS_{L}&\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}&X_{L}\end{array}$
@@
This map $\varrho$ is injective as if we let $L^{\prime}=L^{\mathrm{Ker}(\varrho)},~{}\mathrm{Ker}(\varrho)$ acts trivially on the hexagon, so all $(-1)-$curves are fixed by the kernel, hence they are all defined over $L^{\prime}$ which gives $L^{\prime}=L$, and then $\mathrm{Ker}(\varrho)=\operatorname{Gal}(L/L)$.
The group $\operatorname{Aut}(S_{L})$ (resp. $\operatorname{Aut}(X_{L})$) acts by symmetries on the hexagon of $S_{L}$ (resp.$X_{L}$), which induces homomorphisms
@@
We define $\Delta_{1}$ to be the image of $H$ in $D_{6}$, as $\operatorname{Gal}(L/\mathbb{Q})$ acts on the hexagon faithfully and $\operatorname{Gal}(L/\mathbb{Q})\subseteq D_{6}$, so every element in $\Delta_{1}$ should commute with the image $\tilde{g}$. Since $H\cap(L^{*})^{2}$ is a finite subgroup of $(L^{*})^{2}$, it contains only roots of unity. We define $n$ to be the largest order of a root of unity contained in $H\cap(L^{*})^{2}$. So we have $\mathbb{Q}[\zeta_{n}]$ as an intermediate field of the Galois field extension $L/\mathbb{Q}$, where $\zeta_{n}$ is the primitive $nth$ root of unity. By the fundamental theorem of Galois theory, $L/\mathbb{Q}[\zeta_{n}]$ is Galois, hence we have a surjective group homomorphism $\operatorname{Gal}(L/\mathbb{Q})\rightarrow\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})$, $g\mapsto g\mid_{\mathbb{Q}[\zeta_{n}]}$, this leads to the existence of the following the exact sequence
(4) $1\rightarrow\operatorname{Gal}(L/\mathbb{Q}[\zeta_{n}])\rightarrow\operatorname{Gal}(L/\mathbb{Q})\rightarrow\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})\cong\mathbb{Z}^{*}_{n}\rightarrow 1$
@@
where $0\leq i\leq 5$ and $1\leq j\leq 6$. Then, $H\cong(\mathbb{Z}/6\mathbb{Z})^{2}\rtimes D_{6}$ and thus has $432$ elements. To prove that the surface is rational, the point $P=[1:1:1][1:1:1]$ is fixed by $\tilde{g}$. Let $X_{L}$ be the hexagon over $L$ which is the blow-up of the three points $p_{1}=[1:0:0],~{}p_{2}=[0:1:0]$ and $p_{3}=[0:0:1]$ in $\mathbb{P}^{2}$. Since we have the point $p_{4}=[1:1:1]$ in the middle defined over $L$, then we blow up this point. We will get $X^{\prime}_{L}$ as a Del Pezzo surface of degree $5$, the blow up of $S$ at a $\mathbb{Q}-$rational point is then a Del Pezzo surface of degree $5$ with a $\mathbb{Q}-$rational point and is thus rational [Man66, Theorem 3.15]. ∎
@@
Let $L/\mathbb{Q}$ be the smallest finite Galois extension such that all $(-1)-$curves of $S$ are defined over $L$ and $\mathbb{Q}[\zeta_{n}]$ be an intermediate field, where $n$ is the largest order of a root of unity contained in $L^{*}$ and $\zeta_{n}$ is a primitive $nth$ root of unity. Let $g\in\operatorname{Gal}(L/\mathbb{Q})$ and $\tilde{g}$ be the corresponding Galois action. In order to calculate the elements in $H\cap\left(L^{*}\right)^{2}$, we take $\lambda,\mu\in L^{*}$ and $\tau\in H\cap\left(L^{*}\right)^{2}$, then one can assume that
@@
and $\tau$ commutes with $\tilde{g},~{}i.e.~{}\tau=\tilde{g}^{-1}\tau\tilde{g}$. Let $\operatorname{Gal}(L/\mathbb{Q})$ be isomorphic to a subgroup of $D_{6}$. We apply the exact sequence (4), the group ($\mathbb{Z}_{n})^{*}$ is a quotient of a subgroup of $D_{6}$. The extension $\mathbb{Q}[\zeta_{n}]/\mathbb{Q}$ is abelian and the group $\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})$ is quotient of $\operatorname{Gal}(L/\mathbb{Q})$, $\mathbb{Z}_{n}^{*}$ is abelian. Therefore, it has order at most $m\leq 6$; as $D_{6}$ is not abelian. Hence, using Euler’s totient function, we find that $m=\phi(n)$, which implies that $n$ is at most $18$.
Let $D_{6}=\operatorname{Sym}_{3}\times Z_{2}$, where $\operatorname{Sym}_{3}$ permutes the coordinates $x_{i}$ and $y_{i}$ in the same way, i.e., it is corresponding to
$[x_{1}:x_{2}:x_{3}][y_{1}:y_{2}:y_{3}]\mapsto[x_{\delta(1)}:x_{\delta(2)}:x_{\delta(3)}][y_{\delta(1)}:y_{\delta(2)}:y_{\delta(3)}],$
@@
    $[x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:g^{-1}(\mu)x_{1}:g^{-1}(\lambda)x_{2}][y_{0}:\frac{1}{g^{-1}(\mu)}y_{1}:\frac{1}{g^{-1}(\lambda)}y_{2}]$
@@
    $[x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:\frac{1}{g^{-1}(\mu)}x_{1}:\frac{1}{g^{-1}(\lambda)}x_{2}][y_{0}:g^{-1}(\mu)y_{1}:g^{-1}(\lambda)y_{2}]),$
@@
    If $\operatorname{Gal}(L/\mathbb{Q})$ contains an element that is a $3-$cycle, it corresponds to
@@
    $[x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:g^{-1}(\mu)x_{1}:\frac{g^{-1}(\mu)}{g^{-1}(\lambda)}x_{2}][y_{0}:\frac{1}{g^{-1}(\mu)}y_{1}:\frac{g^{-1}(\lambda)}{g^{-1}(\mu)}y_{2}].$
@@
The remaining cases now do not have any transposition nor any $3-$cycle. It only remains the cases where the group $\operatorname{Gal}(L/\mathbb{Q})$ is trivial or of order $2$. If it is trivial, then $n\leq 2$, so $\mid H\cap\left(L^{*}\right)^{2}\mid\leq 2^{2}$. Hence, $\mid H\mid=\mid H\cap\left(L^{*}\right)^{2}\mid\cdot\mid\Delta_{1}\mid\leq 2^{2}\cdot 12=48$. If $\operatorname{Gal}(L/\mathbb{Q})$ is of order $2$, then $n\leq 6$, so $\mid H\cap\left(L^{*}\right)^{2}\mid\leq 6^{2}$. Hence, $\mid H\mid=\mid H\cap\left(L^{*}\right)^{2}\mid\cdot\mid\Delta_{1}\mid\leq 6^{2}\cdot 12=432$. ∎
5. Del Pezzo surfaces of degree $5$
@@
Let $S$ be the blow-up of the four points $A_{1},..A_{4}$ of $\mathbb{P}^{2}(k)$, in general position. Let $\pi:S\rightarrow\mathbb{P}^{2}$ be the blow-down, then the exceptional divisors on the surface $S$ are given by:
@@
Furthermore, each divisor intersects three others. By [Bla09, Proposition $5.1$], There are thus $5$ sets of $4$ skew exceptional divisors on the surface $S$, namely $\{E_{1},l_{23},l_{24},l_{34}\}$, $\{E_{2},l_{13},l_{14},l_{34}\}$, $\{E_{3},l_{12},l_{14},l_{24}\}$, $\{E_{4},l_{23},l_{13},l_{23}\}$, $\{E_{1},E_{2},E_{3},E_{4}\}$, and the action on these sets gives rise to an isomorphism $\rho:\operatorname{Aut}(S_{k})\rightarrow\operatorname{Sym}_{5}$. ∎
@@
Let $S$ be a Del Pezzo surface of degree $5$ defined over $\mathbb{Q}$. So it is the blow-up of the four points $A_{1},..,A_{4}$ of $\mathbb{P}^{2}(\mathbb{Q})$, in general position. Since under the action of linear automorphisms, all sets of four non-collinear points of $\mathbb{P}^{2}$ are equivalent, we may assume that the four points $A_{1}=(1:0:0),~{}A_{2}=(0:1:0),~{}A_{3}=(0:0:1)$ and $A_{4}=(1:1:1)$ and they are all defined over $\mathbb{Q}$. Furthermore, by Lemma 5.1, $\operatorname{Aut}(S_{\mathbb{Q}})\cong\operatorname{Sym}_{5}$.
@@ 6. Finite subgroups of $\operatorname{PGL}_{3}(\mathbb{Q})$
[Min87, Minkowski’s theorem], [Ser07, Theorem 1]
@@
Since $M(3,2)=4,~{}M(3,3)=1$ and $M(3,p)=0$ for all primes $p>3$, we have
@@
Let $\alpha\in\operatorname{PGL}_{2}(\mathbb{Q})$ be an element of order $3$, By Lemma 6.4, there exists an element $A$ in $\mathrm{SL}_{2}(\mathbb{Q})$ of order $3$ whose class is $\alpha$. By Remark 6.5, $A$ is conjugate to an element $B$ in $\mathrm{SL}_{2}(\mathbb{Z})$ of whose class $\beta$ has order $3$. The quotient of the group $\mathrm{SL}_{2}(\mathbb{Z})$ by its center {±$\{\pm id\}$ is a free product of $\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}/3\mathbb{Z}$ generated by the classes $[S]$ of $S$ and $[T]$ of $T$, where
@@
see [New72, Chap. 8]. Since the class of $S$ has order $3$, then $A$ or $A^{-1}$ is conjugate to the matrix $S$. ∎
@@
We prove that $p\leq 3$ by contradiction. Let $p$ be a prime number $>3$. Let $\alpha\in G$ be an element of order $p$. Since $(p,3)=1$, from Lemma 6.4, there exists $A$ in $\mathrm{SL}_{3}(\mathbb{Q})$ of order $p$ where $\alpha$ is the class of $A$. This implies that the group $H=\langle A\rangle$ in $\mathrm{SL}_{3}(\mathbb{Q})$ has size $p$. By Remark 6.2, $p$ divides $24$, so $p\in\{2,3\}$ which is a contradiction. ∎
@@
If $H$ is a $2-$group in $\operatorname{PGL}_{3}(\mathbb{Q})$, then by Remark 6.2 and Lemma 6.4, we have $\mid H\mid\leq 2^{3}$.
@@
    implies to $u=0,~{}v=0$ and the class of $C$ in $\operatorname{PGL}_{2}(\mathbb{Q})$ is of order $3$, then by Lemma 6.6, the class of $C$ is conjugate to the matrix $\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ or its inverse $\left(\begin{array}[]{cc}-1&1\\ -1&0\end{array}\right)$. We may then replace $C$ by a conjugate and assume that $C=\mu\cdot\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ or $C=\mu\cdot\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ ), where $\mu\in\mathbb{Q}^{*}$. As $A^{3}=Id$, we obtain $\mu=1$. Hence
@@
We prove that there is no element of order $9$, and thus there is no element of order $3^{l}$, where $l>1$. Let $r\leq 5$. Let $\alpha$ be an element of $\operatorname{PGL}_{r}(\mathbb{Q})$ of order $9$. There is thus an element $A$ in $\mathrm{GL}_{r}(\mathbb{Q})$ whose class is $\alpha$ and $a\in\mathbb{Q}$ such that $A^{9}=a\cdot Id$. If we multiply $A$ by a suitable integer number that makes the entries of $A$ integers, we can assume that $a\in\mathbb{Z}$. The polynomial $X^{9}-a$ is then zero when we replace $X$ with $A$ and is thus a multiple of the minimal polynomial of $A$, which is of degree at most $5$. The minimal polynomial of $A$ has then to divide the polynomial $X^{9}-a$. If $a$ is not a cube in $\mathbb{Z}$, then $X^{9}-a$ is irreducible over $\mathbb{Z}$ [Lan02, Theorem 9.1]. So we can we write $a=b^{3}$ where $b$ in $\mathbb{Z}\setminus\{0\}$. This implies that $(X^{9}-b^{3})=(X^{3}-b)(x^{6}+bx^{3}+b^{2})$. We observe that the minimal polynomial cannot divide $X^{3}-b$. Otherwise, we would have $A^{3}=b\cdot Id$, which implies that $\alpha$ is of order $3$. We now look at the polynomial $x^{6}+bx^{3}+b^{2}$. If this one is irreducible, we find a contradiction. In [HJ17, Theorem 3.1], we find that this latter is reducible only when $b^{2}=n^{3}$ and $b=m^{3}-3mn$ for some $m,n$ in $\mathbb{Q}$. We then write $s=b/n$ in $\mathbb{Q}\setminus\{0\}$, and obtain $s^{3}=b^{3}/n^{3}=b^{3}/b^{2}=b$. We then find $a=b^{3}=s^{9}$. The matrix $A^{\prime}=A/s$ is then such that $(A^{\prime})^{9}=Id$ and then we find an element of order $9$ in $\mathrm{GL}_{r}(\mathbb{Q})$. This is impossible: The minimal polynomial should divide $X^{9}-1=(X-1)(X^{2}+X+1)(X^{6}+X^{3}+1)$. As these $3$ are irreducible, it divides $(X-1)(X^{2}+X+1)$ and thus $A^{3}=1$. ∎
@@
Let us assume, for contradiction, the existence of $G\subset\operatorname{PGL}_{3}(\mathbb{Q})$ a group of order $3^{l}$ where $l>1$. Since $G$ is a $3-$group, from [Gri07, Proposition 3.9], the center $G$ of the group $G$ is not trivial and by Lemma 6.10, except the identity, every element in $Z(G)$ is of order $3$. By Lemma 6.9, we can assume that $Z(G)$ contains the class $\beta$ of $B$, where
@@
Let $G\subset\operatorname{PGL}_{3}(\mathbb{Q})$ be a subgroup of finite order $n>1$. let $p$ be a prime number divides $n$, then by Lemma 6.7, $p\leq 3$. So we can write $n=2^{r_{1}}\cdot 3^{r_{2}}$. By Remark 6.8, $r_{1}\leq 3$. By Lemma 6.11, $r_{2}\leq 1$. This implies that $n$ divides $242^{3}\cdot 3^{1}=$ 24. ∎
@@
Let $G$ be the finite group in $\operatorname{PGL}_{3}(\mathbb{Q})$ generated by the following four elements,
@@
7. Del Pezzo surfaces of degree $4$

### v2

@@ 4. Del Pezzo surfaces of degree 66
Over the algebraic closure of $\mathbb{Q}$, we have an isomorphism $S_{\overline{\mathbb{Q}}}\rightarrow X_{\overline{\mathbb{Q}}}$. Let $L/\mathbb{Q}$ be the smallest Galois field extension such that the six (1)(-1)-curves are defined over $L$. The surface $S$ has six (1)(-1)-curves defined over $L$ but not necessarily over $\mathbb{Q}$. The conjugation of the action gives an action of $\operatorname{Gal}(L/\mathbb{Q})$ on $S_{L}$. Indeed, if we write the isomorphism $\varphi:S_{L}\rightarrow X_{L}$ and let $g\in\operatorname{Gal}(L/\mathbb{Q})$, then the following diagram
$\begin{array}[]{ccc}\hskip 14.22636ptS_{L}&\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}&X_{L}\\ g\downarrow&&\downarrow\tilde{g}\\ \hskip 14.22636ptS_{L}&\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}&X_{L}\end{array}$
@@
This map $\varrho$ is injective as if we let $L^{\prime}=L^{\mathrm{Ker}(\varrho)},~\mathrm{Ker}(\varrho)$ acts trivially on the hexagon, so all (1)(-1)-curves are fixed by the kernel, hence they are all defined over $L^{\prime}$ which gives $L^{\prime}=L$, and then $\mathrm{Ker}(\varrho)=\operatorname{Gal}(L/L)$.
[…]
@@
We define $\Delta_{1}$ to be the image of $H$ in $D_{6}$, as $\operatorname{Gal}(L/\mathbb{Q})$ acts on the hexagon faithfully and $\operatorname{Gal}(L/\mathbb{Q})\subseteq D_{6}$, so every element in $\Delta_{1}$ should commute with the image $\tilde{g}$. Since $H\cap(L^{*})^{2}$ is a finite subgroup of $(L^{*})^{2}$, it contains only roots of unity. We define $n$ to be the largest order of a root of unity contained in $H\cap(L^{*})^{2}$. So we have $\mathbb{Q}[\zeta_{n}]$ as an intermediate field of the Galois field extension $L/\mathbb{Q}$, where $\zeta_{n}$ is the primitive nthnth root of unity. By the fundamental theorem of Galois theory, $L/\mathbb{Q}[\zeta_{n}]$ is Galois, hence we have a surjective group homomorphism $\operatorname{Gal}(L/\mathbb{Q})\rightarrow\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})$, $g\mapsto g\mid_{\mathbb{Q}[\zeta_{n}]}$, this leads to the existence of the following the exact sequence
(4) $1\rightarrow\operatorname{Gal}(L/\mathbb{Q}[\zeta_{n}])\rightarrow\operatorname{Gal}(L/\mathbb{Q})\rightarrow\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})\cong\mathbb{Z}^{*}_{n}\rightarrow 1.$
@@
where $0\leq i\leq 5$ and $1\leq j\leq 6$. Then, $H\cong(\mathbb{Z}/6\mathbb{Z})^{2}\rtimes D_{6}$ and thus has $432$ elements. To prove that the surface is rational, the point $P=[1:1:1][1:1:1]$ is fixed by $\tilde{g}$. Let $X_{L}$ be the hexagon over $L$ which is the blow-up of the three points $p_{1}=[1:0:0],~p_{2}=[0:1:0]$ and $p_{3}=[0:0:1]$ in $\mathbb{P}^{2}$. Since we have the point $p_{4}=[1:1:1]$ in the middle defined over $L$, then we blow up this point. We will get $X^{\prime}_{L}$ as a Del Pezzo surface of degree $5$, the blow up of $S$ at a $\mathbb{Q}$-rational point is then a Del Pezzo surface of degree $5$ with a $\mathbb{Q}$-rational point and is thus rational [17, Theorem 3.15]. ∎
@@
Let $L/\mathbb{Q}$ be the smallest finite Galois extension such that all (1)(-1)-curves of $S$ are defined over $L$ and $\mathbb{Q}[\zeta_{n}]$ be an intermediate field, where $n$ is the largest order of a root of unity contained in $L^{*}$ and $\zeta_{n}$ is a primitive nthnth root of unity. Let $g\in\operatorname{Gal}(L/\mathbb{Q})$ and $\tilde{g}$ be the corresponding Galois action. In order to calculate the elements in $H\cap\left(L^{*}\right)^{2}$, we take $\lambda,\mu\in L^{*}$ and $\tau\in H\cap\left(L^{*}\right)^{2}$, then one can assume that
@@
and $\tau$ commutes with $\tilde{g},~i.e.~\tau=\tilde{g}^{-1}\tau\tilde{g}$. Let $\operatorname{Gal}(L/\mathbb{Q})$ be isomorphic to a subgroup of $D_{6}$. We apply the exact sequence (4), the group ($\mathbb{Z}_{n})^{*}$ is a quotient of a subgroup of $D_{6}$. The extension $\mathbb{Q}[\zeta_{n}]/\mathbb{Q}$ is abelian and the group $\operatorname{Gal}(\mathbb{Q}[\zeta_{n}]/\mathbb{Q})$ is quotient of $\operatorname{Gal}(L/\mathbb{Q})$, $\mathbb{Z}_{n}^{*}$ is abelian. Therefore, it has order at most $m\leq 6$; as $D_{6}$ is not abelian. Hence, $m=\phi(n)$, which implies that $n$ is at most $18$, where Eulers totient function $\phi(n)$.
Let $D_{6}=\operatorname{Sym}_{3}\times\mathbb{Z}_{2}$, where $\operatorname{Sym}_{3}$ permutes the coordinates $x_{i}$ and $y_{i}$ in the same way, i.e., it is corresponding to
$x_{1}:x_{2}:x_{3}][y_{1}:y_{2}:y_{3}]\mapsto[x_{\delta(1)}:x_{\delta(2)}:x_{\delta(3)}][y_{\delta(1)}:y_{\delta(2)}:y_{\delta(3)}],$
@@
    $x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:g^{-1}(\mu)x_{1}:g^{-1}(\lambda)x_{2}][y_{0}:\frac{1}{g^{-1}(\mu)}y_{1}:\frac{1}{g^{-1}(\lambda)}y_{2}]$
@@
    $x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:\frac{1}{g^{-1}(\mu)}x_{1}:\frac{1}{g^{-1}(\lambda)}x_{2}][y_{0}:g^{-1}(\mu)y_{1}:g^{-1}(\lambda)y_{2}]),$
@@
    If $\operatorname{Gal}(L/\mathbb{Q})$ contains an element that is a 33-cycle, it corresponds to
@@
    $x_{0}:x_{1}:x_{2}][y_{0}:y_{1}:y_{2}]\mapsto[x_{0}:g^{-1}(\mu)x_{1}:\frac{g^{-1}(\mu)}{g^{-1}(\lambda)}x_{2}][y_{0}:\frac{1}{g^{-1}(\mu)}y_{1}:\frac{g^{-1}(\lambda)}{g^{-1}(\mu)}y_{2}].$
@@
The remaining cases now do not have any transposition nor any 33-cycle. It only remains the cases where the group $\operatorname{Gal}(L/\mathbb{Q})$ is trivial or of order $2$. If it is trivial, then $n\leq 2$, so $\mid H\cap\left(L^{*}\right)^{2}\mid\leq 2^{2}$. Hence, $\mid H\mid=\mid H\cap\left(L^{*}\right)^{2}\mid\cdot\mid\Delta_{1}\mid\leq 2^{2}\cdot 12=48$. If $\operatorname{Gal}(L/\mathbb{Q})$ is of order $2$, then $n\leq 6$, so $\mid H\cap\left(L^{*}\right)^{2}\mid\leq 6^{2}$. Hence, $\mid H\mid=\mid H\cap\left(L^{*}\right)^{2}\mid\cdot\mid\Delta_{1}\mid\leq 6^{2}\cdot 12=432$. ∎
5. Del Pezzo surfaces of degree 55
@@
Let $S$ be the blow-up of the four points $A_{1},..,A_{4}$ of $\mathbb{P}^{2}(k)$, in general position. Let $\pi:S\rightarrow\mathbb{P}^{2}$ be the blow-down, then the exceptional divisors on the surface $S$ are given by:
@@
Furthermore, each divisor intersects three others. By [5, Proposition $5.1$], There are thus $5$ sets of $4$ skew exceptional divisors on the surface $S$, namely $\{E_{1},l_{23},l_{24},l_{34}\}$, $\{E_{2},l_{13},l_{14},l_{34}\}$, $\{E_{3},l_{12},l_{14},l_{24}\}$, $\{E_{4},l_{23},l_{13},l_{23}\}$, $\{E_{1},E_{2},E_{3},E_{4}\}$, and the action on these sets gives rise to an isomorphism $\rho:\operatorname{Aut}(S_{k})\rightarrow\operatorname{Sym}_{5}$. ∎
@@
Let $S$ be the Del Pezzo surface of degree $5$ defined over $\mathbb{Q}$ that is the blow-up of the four points $A_{1},..,A_{4}$ of $\mathbb{P}^{2}(\mathbb{Q})$, in general position. Since under the action of linear automorphisms, all sets of four non-collinear points of $\mathbb{P}^{2}$ are equivalent, we may assume that the four points $A_{1}=(1:0:0),~A_{2}=(0:1:0),~A_{3}=(0:0:1)$ and $A_{4}=(1:1:1)$ and they are all defined over $\mathbb{Q}$. Furthermore, by Lemma 5.1, $\operatorname{Aut}(S_{\mathbb{Q}})\cong\operatorname{Sym}_{5}$.
@@ 6. Finite subgroups of $\operatorname{PGL}_{3}(\mathbb{Q})$
[18, Minkowskis theorem], [21, Theorem 1]
@@
Since M(3,2)=4,M(3,3)=1M(3,2)=4,~M(3,3)=1 and $M(3,p)=0$ for all primes $p>3$, we have
@@
Let $\alpha\in\operatorname{PGL}_{2}(\mathbb{Q})$ be an element of order $3$, By Lemma 6.4, there exists an element $A$ in $\mathrm{SL}_{2}(\mathbb{Q})$ of order $3$ whose class is $\alpha$. By Remark 6.5, $A$ is conjugate to an element $B$ in $\mathrm{SL}_{2}(\mathbb{Z})$ of whose class $\beta$ has order $3$. The quotient of the group $\mathrm{SL}_{2}(\mathbb{Z})$ by its center $\{\pm id\}$ is a free product of $\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}/3\mathbb{Z}$ generated by the classes $[S]$ of $S$ and $[T]$ of $T$, where
@@
see [19, Chap. 8]. Since the class of $S$ has order $3$, then $A$ or $A^{-1}$ is conjugate to the matrix $S$. ∎
@@
We prove that $p\leq 3$ by contradiction. Let $p$ be a prime number >3>3. Let $\alpha\in G$ be an element of order $p$. Since $(p,3)=1$, from Lemma 6.4, there exists $A$ in $\mathrm{SL}_{3}(\mathbb{Q})$ of order $p$ where $\alpha$ is the class of $A$. This implies that the group $H=\langle A\rangle$ in $\mathrm{SL}_{3}(\mathbb{Q})$ has size pp. By Remark 6.2, $p$ divides $24$, so $p\in\{2,3\}$ which is a contradiction. ∎
@@
If $H$ is a 22-group in $\operatorname{PGL}_{3}(\mathbb{Q})$, then by Remark 6.2 and Lemma 6.4, we have $\mid H\mid\leq 2^{3}$.
@@
    implies to u=0,v=0u=0,~v=0 and the class of $C$ in $\operatorname{PGL}_{2}(\mathbb{Q})$ is of order $3$, then by Lemma 6.6, the class of $C$ is conjugate to the matrix $\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ or its inverse $\left(\begin{array}[]{cc}-1&1\\ -1&0\end{array}\right)$. We may then replace $C$ by a conjugate and assume that $C=\mu\cdot\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ or $C=\mu\cdot\left(\begin{matrix}0&-1\\ 1&-1\end{matrix}\right)$ ), where $\mu\in\mathbb{Q}^{*}$. As $A^{3}=Id$, we obtain $\mu=1$. Hence
@@
We prove that there is no element of order $9$, and thus there is no element of order $3^{l}$, where $l>1$. Let $r\leq 5$. Let $\alpha$ be an element of $\operatorname{PGL}_{r}(\mathbb{Q})$ of order $9$. There is thus an element $A$ in $\mathrm{GL}_{r}(\mathbb{Q})$ whose class is $\alpha$ and $a\in\mathbb{Q}$ such that $A^{9}=a\cdot Id$. If we multiply $A$ by a suitable integer number that makes the entries of $A$ integers, we can assume that $a\in\mathbb{Z}$. The polynomial $X^{9}-a$ is then zero when we replace $X$ with $A$ and is thus a multiple of the minimal polynomial of $A$, which is of degree at most $5$. The minimal polynomial of $A$ has then to divide the polynomial $X^{9}-a$. If $a$ is not a cube in $\mathbb{Z}$, then $X^{9}-a$ is irreducible over $\mathbb{Z}$ [16, Theorem 9.1]. So we can we write $a=b^{3}$ where $b$ in $\mathbb{Z}\setminus\{0\}$. This implies that $(X^{9}-b^{3})=(X^{3}-b)(x^{6}+bx^{3}+b^{2})$. We observe that the minimal polynomial cannot divide $X^{3}-b$. Otherwise, we would have $A^{3}=b\cdot Id$, which implies that $\alpha$ is of order $3$. We now look at the polynomial $x^{6}+bx^{3}+b^{2}$. If this one is irreducible, we find a contradiction. In [11, Theorem 3.1], we find that this latter is reducible only when $b^{2}=n^{3}$ and $b=m^{3}-3mn$ for some $m,n$ in $\mathbb{Q}$. We then write $s=b/n$ in $\mathbb{Q}\setminus\{0\}$, and obtain $s^{3}=b^{3}/n^{3}=b^{3}/b^{2}=b$. We then find $a=b^{3}=s^{9}$. The matrix $A^{\prime}=A/s$ is then such that $(A^{\prime})^{9}=Id$ and then we find an element of order $9$ in $\mathrm{GL}_{r}(\mathbb{Q})$. This is impossible: The minimal polynomial should divide $X^{9}-1=(X-1)(X^{2}+X+1)(X^{6}+X^{3}+1)$. As these $3$ are irreducible, it divides $(X-1)(X^{2}+X+1)$ and thus $A^{3}=1$. ∎
@@
Let us assume, for contradiction, the existence of $G\subset\operatorname{PGL}_{3}(\mathbb{Q})$ a group of order $3^{l}$ where $l>1$. Since $G$ is a 33-group, from [10, Proposition 3.9], the center $G$ of the group $G$ is not trivial and by Lemma 6.10, except the identity, every element in $Z(G)$ is of order $3$. By Lemma 6.9, we can assume that $Z(G)$ contains the class $\beta$ of $B$, where
@@
Let $G\subset\operatorname{PGL}_{3}(\mathbb{Q})$ be a subgroup of finite order $n>1$. let $p$ be a prime number divides $n$, then by Lemma 6.7, $p\leq 3$. So we can write $n=2^{r_{1}}\cdot 3^{r_{2}}$. By Remark 6.8, $r_{1}\leq 3$. By Lemma 6.11, $r_{2}\leq 1$. This implies that $n$ divides $2^{3}\cdot 3^{1}=24$. ∎
@@
Let $G$ be the finite group in $\operatorname{PGL}_{3}(\mathbb{Q})$ generated by the following three elements,
@@
7. Del Pezzo surfaces of degree 44